The sum of two numbers is 39. Their difference is 7. What are the two numbers? Show both equations and your method for solving as well as the solution.

Answer by ptaylor(2198)

Let x=one number
And y=the other
Now we are told the following:
x+y=39------------------eq1
x-y=7----------------------eq2
Add eq1 and eq2
2x=46 divide both sides by 2
x=23 one number
substitute x=23 into eq1
23+y=39 subtract 23 from both sides
23-23+y=39-23 collect like terms
y=16 the other number
ANOTHER WAY:
Let x=one number
Then 39-x=the other number

Now we are told that:
(39-x)-x=7
39-2x=7 subtract 39 from both sides
-2x=7-39 collect like terms
-2x=-32 divide both sides by -2
x=16---one number
39-x=39-16=23----------the other number
CK
16+23=39
39=39
and
23-16=7
7=7

Hope this helps----ptaylor